OpenStudy (anonymous):

What is the value of -2^-2^-2 ?

OpenStudy (rajat97):

-2^-2 can be written as (1/-2)^2=1/4 and (1/4)^-2=4^2 if the power is -ve, we take the reciprocal of the number and raise it to the magnitude of the power(the -ve power becomes +ve)

OpenStudy (rajat97):

and 4^2=16

OpenStudy (rajat97):

and 16 is the answer